OpenStudy (anonymous):

Find the general solution \[\left( {2 \over \sqrt {1 - x^2}} + y \cos(xy)\right)dx + \left( {x \cos(xy) - y^{-{1 \over 3}}} \right)dy= 0\]

OpenStudy (turingtest):

this is exact, eh?

OpenStudy (anonymous):

It is!

OpenStudy (turingtest):

\[\left( {2 \over \sqrt {1 - x^2}} + y \cos(xy)\right) + \left( {x \cos(xy) - y^{-{1 \over 3}}} \right)\frac{dy}{dx}= 0\]\[\int x\cos(xy)-y^{-1/3}dy=\sin(xy)-\frac32y^{2/3}+h(x)\]something like that... somebody jump in, it's been a while for me with these

OpenStudy (anonymous):

I think that i have this one, it's just going to take me a while to work it all out.

OpenStudy (anonymous):

My brain is fried, I've been working on these too long today

OpenStudy (turingtest):

yeah take the partial wrt x on this guy and compare it to the other term blah, blah... me too, I'm up to my head in python... programming's not my thing

OpenStudy (anonymous):

Python is cool; I almost cried in despair writing a C project (granted, I did badly with the data abstraction and interfaces). And this should be messy but not that complicated. If you want, post your work @gmer after you are done and I can do a check for you.

OpenStudy (turingtest):

\[\Psi_x=y\cos(xy)-y^{-1/3}+h'(x)\]so\[h'(x)={2\over\sqrt{1-x^2}}\]which integrates to some arcsin of something I think... @bmp I'm so far behind in comp sci, I feel like I'm learning the alphabet print 'hello world' lol

OpenStudy (anonymous):

I've already almost forgotten what python looks like after not touching it for a year. I should probably keep up on my c and java before it happens with them too :)

OpenStudy (anonymous):

Is this c++?

OpenStudy (anonymous):

@TuringTest My project is still broken, so it happens haha. I almost came up with a (very bad) Sudoku solver using backtracking, but I couldn't :-(. The only cool thing about my big blob of code (honestly) is that I could do the animation using only ANSI escape sequences on a terminal. Not really dynamic, but it was cool. @gmer I only remember basic Python stuff too; I use it to parse data on my research lab and that's all. I wanted to learn about the (now deprecated) urllib, sockets, and so on.

OpenStudy (turingtest):

may I ask your field of research @bmp ?

OpenStudy (anonymous):

@TuringTest I started working what people tend to call condensed physics, photoacoustic spectroscopy. I will begin working on Raman scattering and spectroscopy, if everything goes well. But as an undergrad, I am still taking baby steps on the lab. :-)

OpenStudy (turingtest):

Oh how I envy those with lab access if you're taking baby steps I must be a fetus

OpenStudy (anonymous):

Yeah, the lab is cool. But I thought you were a major (or majoring) in the Maths. :-)

OpenStudy (turingtest):

nope, self-taught

OpenStudy (anonymous):

Ah, that makes us two. It's interesting to note that this seems like a trend with the resources avaiable on the internet. But maybe I am just rambling now :-)

OpenStudy (turingtest):

Not to me. I see resources like this and OCW as vast resources for people like you and I. Long live the infinite library of the internet!